OpenStudy (loser66):

Suppose g:R→R is an everywhere differentiable function. Show that if g(a)≠0, then |g(x)| is differentiable at x =a. Hint: use the fact that composition of 2 differentiable functions is differentiable. b) Now suppose g(a) =0. Show that |g(x)| is differentiable at x = a iff g'(a) =0 Please, help

OpenStudy (anonymous):

i'll try tomorrow

OpenStudy (loser66):

ok!! good night, friend.

OpenStudy (loser66):

This is the second time I post this problem. It seems no one wants to help me. :(

OpenStudy (anonymous):

not like that @Loser66 few ppl here only know these stuff dont worry :)

OpenStudy (kainui):

I don't know how to do this, I am just here watching sorry.

OpenStudy (loser66):

@Kainui are you kidding me? it 's calculus.

OpenStudy (loser66):

and I know you are an Ace on the field

OpenStudy (kainui):

That's different, I don't believe in proofs.

OpenStudy (anonymous):

lol

OpenStudy (loser66):

hahaha..... it's ok!!

OpenStudy (anonymous):

part 1

OpenStudy (anonymous):

that it only

OpenStudy (anonymous):

part 2, g(a)=0 means we need to check defferiantability at g(a)=0 when g'(a)=0 means its smooth so its deferentiable

OpenStudy (anonymous):

i wrote these hints hope it helps you to write neat answer

OpenStudy (loser66):

Suppose g(a) =0, then g is differentiable at x =a \(g'(a) = lim_{x\rightarrow a}\dfrac{g(x)- g(a)}{x-a}= lim_{x\rightarrow a}\dfrac{g(x)}{x-a}\) For other piecewise, we have the same argument, and let them equal.

OpenStudy (loser66):

That is \(lim_{x\rightarrow a}\dfrac{g(x)}{x-a} =lim_{x\rightarrow a}\dfrac{-g(x)}{x-a}\) then?? how to argue to let to g'(a) =0??